# Competitive Equilibrium for Chores: from Dual Eisenberg-Gale to a Fast, Greedy, LP-based Algorithm

### v1

@@ 5. Experiments
In Section 4, we showed that the GFW algorithm guarantees convergence to a strongly approximate CE in polynomial time, and converges to an exact CE in finite number of iterations, while requiring only one LP solve per iteration. In contrast, the prior numerical state-of-the-art method, the EPM of Boodaghians et al. (2022) requires solving a convex quadratic program (QP) at every iteration. We remark that we do not provide comparisons with the combinatorial algorithm ((Chaudhury et al., 2022)). This is primarily attributed to the numerical issues that arise with the arithmetic performed by combinatorial algorithms– in particular, the bit-length of the prices and the allocation can grow exponentially with the iterations in the combinatorial algorithms, as conjectured by (Duan and Mehlhorn, 2015). [^11] In private communication, we are aware that Omar Darwish and Kurt Mehlhorn have an implementation of the combinatorial algorithm for Arrow-Debreu markets with goods, and they confirm the conjecture in (Duan and Mehlhorn, 2015).
@@
In this section, we investigate the numerical performance of GFW as compared to the EPM, when both utilize state-of-art optimization software for solving LPs and QPs (in particular, using Gurobi). In Algorithm 1 we give pseudocode for the EPM for computing an exact CE for chores. The set $\mathcal{D}^{+}$ in the pseudocode denotes the set of all possible disutility profiles corresponding to any feasible allocation, or over-allocation, i.e.,
@@
$\mathcal{D}^{+}=\left\{d\in\mathbb{R}^{n}_{\geq 0}\;|\;\exists\;x\in\mathbb{R}^{n\times m}_{\geq 0}\mbox{ such that }d_{i}\geq\sum\nolimits_{j}d_{ij}x_{ij}\;\forall\;i\in[n],\sum\nolimits_{i}x_{ij}=1\;\forall\;j\in[m]\right\}.$
Input: An initial infeasible disutility profile $d^{0}$ such that $d^{0}_{i}>0$ and close to 00 for all $i\in[n]$
for $k=1,2,\ldots$ do
      if $d^{k}$ is infeasible then
           $d^{k,*}={\arg\min}_{d\in\mathcal{D}^{+}\cap\{d|d\geq d^{k}\}}\big{\lVert}d-d^{k}\big{\rVert}$ 1212footnotemark: 12;
           $a^{k}\leftarrow{(d^{k,*}-d^{k})n}\big{/}{\langle d^{k,*}-d^{k},d^{k,*}\rangle}$;
           $d^{k+1}\leftarrow{1}\big{/}{a^{k}}$;
           else
                return $(d^{k},a^{k-1})$;
@@
We compare the algorithms on a variety of instances. First, we consider randomly generated instances. For a given instance size $n$, we construct $n$ buyers, each with a budget of $1$. Then, we generate $n$ chores. This means that disutilities are an $n\times n$ matrix $D$. Each cell $d_{ij}$ in $D$ is sampled iid. from some distribution $d_{ij}\sim F$. We consider five different distributions of disutilities: uniform on $[0,1)$, log-normal associated with the standard normal distribution $\mathcal{N}(0,1)$, truncated normal associated with $\mathcal{N}(0,1)$, and truncated at ${10}^{-3}$ and $10$ standard deviations from 00, exponential with the scale parameter $1$, and uniform random integers on $\{1,\ldots,1000\}$. For each distribution, we consider instance sizes $n\in\{2,50,100,150,200,250,300\}$, and for each choice of distribution and $n$, we generate 100 instances.
@@
Secondly, we generate a set of chores instances by considering a dataset of bids from PC members at the AAMAS conference on potential papers to review, obtained from PrefLib (Mattei and Walsh, 2013). We use the 2021 dataset, which has 596 PC members included in the dataset. Each PC member scored each paper ordinally as {yes ¿ maybe ¿ no response ¿ no ¿ conflict}. In order to generate real-valued disutilities, we convert these into the values {1, 3, 5, 7, $4000$}. We then generate two different datasets: In the first, we use these 5 values directly, and sample subsets of PC members and papers as follows: first uniformly choose one paper at random, then choose a subset of papers which are the most similar to it (considering the PC members’ responses as feature vectors of the paper), finally choose the subset of PC members that have the largest number of non-conflict responses on the chosen papers. In the second, we additionally add Gaussian noise to each valuation, using $\mathcal{N}(0,0.04)$, and generate 100 instances.
We ran all experiments on a personal desktop that uses the Apple M1 Chip and has 8GB RAM. We used Gurobi version 11.0.0.
Since both algorithms are known to be able to find an $\epsilon$-strongly approximate CE in polynomial time, we measure the level of approximation by plotting the corresponding $\epsilon$ for each algorithm. Furthermore, we note that the EPM finds an exact CE when the disutility profile generated for the next iterate is in the feasible set. To compare with the GFW algorithm, we run the EPM until a feasible (or over-allocated) disutility profile is found. In our experiments, we consider any CE with $\epsilon\leq 0.01$ as an approximate CE and any CE with $\epsilon\leq 10^{-6}$ as an exact CE.
@@
To measure the equilibrium approximation quality $\epsilon$ at a given iteration, we compute a pair $(p,x)$ corresponding to the current iterate for each algorithm. For GFW, we compute $(p,x)$ based on Lemma 5, where $x$ is given by the optimal dual variable of the corresponding LP. For EPM, we solve an LP to find a feasible $x$ corresponding to the final $d^{k}$ if the algorithm finds an exact CE, and retrieve $x$ corresponding to $d^{k,*}$ if $(d^{k,*},a^{k})$ corresponds to a strongly approximate CE. We compute prices $p$ for EPM via $p_{j}=\min_{i}a_{i}d_{ij}$ (Boodaghians et al., 2022).
@@
Though the EPM is supposed to reach an exact CE when it finds a feasible (or over-allocated) disutility profile, we found in our experiments that an unignorable numerical error can be caused by the approximate solution of QPs. If the projection step in Algorithm 1 cannot be solved accurately, this can cause $a^{k-1}$, the normal vector of the generated hyperplane, to point in the wrong direction. This leads to an inaccurate price vector $p^{*}$. Moreover, let $d^{k}$ be the true iterate with a perfect-precision QP solution, and let $\tilde{d}^{k}$ be the “solution” found by Gurobi at a problematic iteration. The Gurobi solution $\tilde{d}^{k}$ may end up being feasible even if $d^{k}$ is not, or vice versa. This may lead to incorrect termination decisions in either case (in which case our attempt to compute a corresponding CE allocation fails). This is a critical issue, because we have no way to proceed once this happens. Considering this issue, we say that the EPM fails if we cannot find a corresponding strongly approximate CE even though the EPM ostensibly terminated with one.
Our results on the five different distributions for iid random disutilities are shown in Fig. 1.
@@
Figure 1. Numerical results on iid randomly generated disutility matrix instances. Each row corresponds to a different disutility distribution (noted on the left). Left: Average number of iterations to solve a given instance size. Middle: Average wall clock running time to solve a given instance size (using Gurobi for both algorithms). Right: Fraction of instances solved for each algorithm. GFW solves every instance for every distribution.
@@
Overall, we conclude that GFW is the first highly practical algorithm: it solves every instance in very little time and in few iterations, even for instances where the size of the disutility matrix is $300\times 300$. In comparison, EPM fails to solve many instances in our datasets. As discussed before, we do caution that the success of EPM is highly dependent on the performance of the Gurobi QP solver, and in particular its accuracy. We found that the QP performance of Gurobi can vary substantially across different machines, and unpredictably so. For example, Gurobi may choose different methods to solve LPs or QPs based on how much RAM is available, as well as other factors. Due to this so-called ”performance variability”, it is difficult to conclusively evaluate the wall-clock performance of the algorithms, as both algorithms rely on the performance of Gurobi on large-size instances. In particular, some preliminary experiments suggests that the wall-clock performance of EPM improves relative to the wall-clock performance of GFW if one adds more RAM. However, we found that even in this case, EPM still has a significant failure rate, since Gurobi still outputs approximate QP solutions, and this still leads to numerical errors that break EPM.
@@
Figure 2. Numerical results on AAMAS PC member bidding data. The top row uses the “original” valuations. The bottom row adds Gaussian noise. Left: Average number of iterations to solve a given instance size. Middle: Average wall clock running time to solve a given instance size. Right: Fraction of instances solved.
6. Conclusion and Discussion
We develop a new convex maximization program for computing CE in Fisher markets with chores, and showed that it satisfies some dual-like properties with the concave minimization program for computing CE. We then derived a new redundant constraint which allows us to avoid directions that lead to infinitely positive objective. This yielded the first mathematical program for finding CE for chores such that it is possible to run iterative optimization methods over a convex feasible region. We then introduced the GFW method for computing CE. From prior results on convex maximization over polytopes, it follows that this method terminates in a finite number of iterations. We then showed that, in fact, the method finds an $\epsilon$-approximate CE in $O(1/\epsilon^{2}$ iterations, while requiring only a simple LP solve at every iteration. We also gave a more general convergence result that generalizes prior results for strongly convex functions to Bregman-strongly-convex functions. Finally, we showed through extensive numerical experiments that our method is highly practical: it solved every instance that we tried in a minute or less and is both robust and simple to implement. In contrast, we showed that the prior state-of-the-art method, the EPM, often fails to solve instances altogether, due to numerical accuracy problems that break the algorithm.
Challenge of Moving from Approximate CE to Exact CE:
The obvious open problem is to settle the exact complexity of finding a CE in the chores market. A natural attempt would be to find a good approximate CE and then argue how to move to a “nearby” exact CE. Unfortunately, such approaches can face the issue that there are instances where an approximate CE can be very far from an exact CE (See Section B in Appendix for an example).
@@ 6. Conclusion and Discussion
Despite the above hurdle, we believe that finding a $\varepsilon$-CE in the chores market with a better run time dependence on $1/\varepsilon$ is of theoretical interest. An ideal goal would be to have a poly-logarithmic dependence on $1/\varepsilon$, but any polynomial improvement would be a stepping stone (currently, all methods have a dependence of $\mathcal{O}(1/\varepsilon^{2})$).
Acknowledgments
Christian Kroer was supported by the Office of Naval Research awards N00014-22-1-2530 and N00014-23-1-2374, and the National Science Foundation awards IIS-2147361 and IIS-2238960. Ruta Mehta was supported by NSF grant CCF-2334461 and NSF CAREER Award CCF-1750436.
@@ Appendix A A Primal-Dual Analysis of the Convergence of the Frank Wolfe Method to an Approximate Equilibrium
Since $f$ does not depend on $p$, we ease notation in this by section by writing $f(\beta)$ for the Chores Dual objective function, which was previously defined as depending on both $p$ and $\beta$.
@@
(20) $\min_{\beta}\quad$ $\sum_{i\in[n]}B_{i}\cdot\frac{\beta_{i}}{\beta^{t-1}_{i}}$
@@
[…]
Claim 1 ().
@@
[…]
@@
From Claim 2, it follows that $\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t-1}_{ij})^{B_{i}}=(\prod_{i\in[n]}(1/\beta^{t-1}_{i})^{B_{i}})\cdot(\prod_{i\in[n]}(\sum_{j\in[m]}$ $p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}})$ and Claim 1 it follows that $\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t}_{ij})^{B_{i}}=\prod_{i\in[n]}(B_{i}/\beta^{t-1}_{i})^{B_{i}}$ Therefore, if we comparing the weighted product of disutilities in the current iteration and the previous iteration, we have
$\frac{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t}_{ij})^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t-1}_{ij})^{B_{i}}}$ $=\frac{\prod_{i\in[n]}(B_{i}/\beta^{t-1}_{i})^{B_{i}}}{(\prod_{i\in[n]}(1/\beta^{t-1}_{i})^{B_{i}})\cdot(\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}})}$
@@
[…]
Claim 3 ().
@@
Under stationarity conditions, we have $\sum_{i\in[n]}x^{t}_{ij}=-\delta^{t}$ for all $j$ such that $p^{t}_{j}>0$. Furthermore, from Claim 2, it follows that $\sum_{j\in[m]}p^{t}_{j}x^{t}_{ij}=\beta^{t}_{i}\cdot(\sum_{j\in[m]}d_{ij}x^{t}_{ij})$. Observe that,
@@
$=\sum_{i\in[n]}\beta^{t}_{i}\cdot(\sum_{j\in[m]}d_{ij}x^{t}_{ij})$ (by Claim 1)
$=\sum_{i\in[n]}\sum_{j\in[m]}p^{t}_{j}x^{t}_{ij}$ (by Claim 2)
@@
$=\sum_{j\mid p^{t}_{j}>0}p^{t}_{j}\cdot(-\delta^{t})=(-\delta^{t})\cdot\sum_{j\in[m]}p^{t}_{j}$
@@
[…]
Claim 4 ().
If $(p^{t-1},x^{t-1})$ is not an $\varepsilon$-approximate equilibrium, then $\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}\geq\min_{i}B_{i}\cdot\varepsilon^{2}/2.3$.

### v2

@@ 5 Experiments
In Section 4, we showed that the GFW algorithm guarantees convergence to a strongly approximate CE in polynomial time, and converges to an exact CE in finite number of iterations, while requiring solving only one LP per iteration. In contrast, the prior numerical state-of-the-art method, the EPM of Boodaghians et al. (2022) requires solving a convex quadratic program (QP) at every iteration.
In this section, we investigate the numerical performance of GFW as compared to the EPM, when both utilize state-of-the-art optimization software for solving LPs and QPs (in particular, using Gurobi). We remark that we do not provide comparisons with the combinatorial algorithm (Chaudhury et al. 2022). This is primarily attributed to the numerical issues that arise with the arithmetic performed by combinatorial algorithms–in particular, the bit-length of the prices and the allocation can grow exponentially with the iterations in the combinatorial algorithms, as conjectured by Duan and Mehlhorn (2015). [^6] In private communication, we are aware that Omar Darwish and Kurt Mehlhorn have an implementation of the combinatorial algorithm for Arrow-Debreu markets with goods, and they confirm the conjecture in Duan and Mehlhorn (2015).
@@
In 10 we give pseudocode for the EPM for computing an exact CE for chores. The set $\mathcal{D}^{+}$ in the pseudocode denotes the set of all possible disutility profiles corresponding to any feasible allocation or over-allocation, i.e.,
@@
$\mathcal{D}^{+}=\Big{\{}d\in\mathbb{R}^{n}_{\geq 0}\;|\;\exists\;x\in\mathbb{R}^{n\times m}_{\geq 0}\mbox{ such that }d_{i}\geq\sum\nolimits_{j}d_{ij}x_{ij}\;\forall\;i\in[n],\sum\nolimits_{i}x_{ij}=1\;\forall\;j\in[m]\Big{\}}.$
[…]
2:for $k=1,2,\ldots$ do
3:     if $d^{k}$ is infeasible then
4:         $d^{k,*}={\arg\min}_{d\in\mathcal{D}^{+}\cap\{d|d\geq d^{k}\}}\big{\lVert}d-d^{k}\big{\rVert}$ 77footnotemark: 7
5:         $a^{k}\leftarrow{(d^{k,*}-d^{k})n}\big{/}{\langle d^{k,*}-d^{k},d^{k,*}\rangle}$
6:         $d^{k+1}\leftarrow{1}\big{/}{a^{k}}$
7:     else
8:         return $(d^{k},a^{k-1})$
9:     end if
10:end for
@@
We compare the algorithms on a variety of instances. First, we consider randomly generated instances. For a given instance size $n$, we construct $n$ buyers, each with a budget of $1$. Then, we generate $n$ chores. This means that disutilities are an $n\times n$ matrix $D$. Each cell $d_{ij}$ in $D$ is sampled i.i.d. from some distribution $F$: $d_{ij}\sim F$. We consider five different distributions of disutilities: (1) uniform on $[0,1)$; (2) log-normal associated with the standard normal distribution $\mathcal{N}(0,1)$; (3) truncated normal associated with $\mathcal{N}(0,1)$ and truncated at ${10}^{-3}$ and $10$ standard deviations from 00; (4) exponential with the scale parameter $1$; (5) uniform random integers on $\{1,\ldots,1000\}$. For each distribution, we consider instance sizes $n\in\{2,50,100,150,200,250,300\}$. For each choice of distribution and $n$, we generate 100 instances.
@@
Secondly, we generate a set of chores instances by considering a dataset of bids from PC members at the AAMAS conference on potential papers to review, obtained from PrefLib (Mattei and Walsh 2013). We use the 2021 dataset, which has 596 PC members included in the dataset. Each PC member scored each paper ordinally as {yes >>> maybe >>> no response >>> no >>> conflict}. In order to generate real-valued disutilities, we convert these into the values {1, 3, 5, 7, $4000$}. We then generate two different datasets: In the first, we use these 5 values directly, and sample subsets of $n$ PC members and $n$ papers as follows: first uniformly choose one paper at random, then choose a subset of papers (of size $n$) which are the most similar to it (considering the PC members’ responses as feature vectors of the paper), finally choose the subset (of size $n$) of PC members that have the largest number of non-conflict responses on the chosen papers. In the second, we additionally add Gaussian noise to each valuation, using $\mathcal{N}(0,0.04)$. For both datasets, we consider instance size $n\in\{2,50,100,150,200,250,300\}$. For each dataset and $n$, we generate 100 instances.
We ran all experiments on a personal desktop that uses the Apple M1 Chip and has 8GB RAM. We used Gurobi version 11.0.0. Since the tolerance parameters for Gurobi models may influence the accuracy of the solutions per iteration and thus the performance of EPM, we tuned those parameters. We considered 4 parameters: we tuned FeasibilityTol in the range of $10^{-9}$ to $10^{-6}$, OptimalityTol in the range of $10^{-9}$ to $10^{-6}$, BarConvTol in the range of 00 to $10^{-8}$, and BarCorrectors in the range of 00 to $10^{4}$  [^8] The default setting chooses this parameter automatically, depending on problem characteristics. . For our final experiments, we set these parameters as FeasibilityTol = $10^{-6}$ (default), OptimalityTol = $10^{-6}$ (default), BarConvTol = 00 (minimum), and BarCorrectors as default. This is because shrinking BarConvTol significantly increases the accuracy of solutions; but decreasing the other tolerance parameters, or increasing BarCorrectors, greatly increases the time Gurobi spends per iteration, thus harming the EPM performance, with no significant improvement to the solution accuracy. See Appendix C for plots illustrating this point.
@@
In our experiments, we measure the computational cost required to find a strongly-approximate CE and an exact CE. In particular, we measure the number of iterations and the running time. Recall the definitions of $\epsilon$-approximate CE and exact CE in Section 2. We consider any condition with $\epsilon\leq 10^{-6}$ as a condition that holds exactly. We consider any CE with the first condition relaxed to $\epsilon\leq 0.01$ (and the other two conditions exactly satisfied) as a strongly-approximate CE. To measure the equilibrium approximation quality $\epsilon$ at a given iteration, we compute a pair $(p,x)$ corresponding to the current iterate for each algorithm. For GFW, we compute $(p,x)$ based on Lemma 2, where $x$ is given by the optimal dual variable of the corresponding LP. For EPM, we solve an LP to find a feasible $x$ corresponding to the final $d^{k}$ if the algorithm finds an exact CE, and retrieve $x$ corresponding to $d^{k,*}$ if $(d^{k,*},a^{k})$ corresponds to a strongly-approximate CE. We compute prices $p$ for EPM via $p_{j}=\min_{i}a_{i}d_{ij}$ (Boodaghians et al. 2022).
@@
Though the EPM is supposed to reach an exact CE when it finds a feasible (or over-allocated) disutility profile, we found in our experiments that an unignorable numerical error can be caused by the approximate solution of QPs. If the projection step in 10 cannot be solved accurately, this can cause $a^{k-1}$, the normal vector of the generated hyperplane, to point in the wrong direction. This leads to an inaccurate price vector $p^{*}$. Moreover, let $d^{k}$ be the true iterate with a perfect-precision QP solution, and let $\tilde{d}^{k}$ be the “solution” found by Gurobi at a problematic iteration. The Gurobi solution $\tilde{d}^{k}$ may end up being feasible even if $d^{k}$ is not, or vice versa. This may lead to incorrect termination decisions in either case (in which case our attempt to compute a corresponding CE allocation fails). This is a critical issue, because we have no way to proceed once this happens. Considering this issue, we say that the EPM fails if we cannot find a corresponding strongly approximate CE even though the EPM ostensibly terminated with one. In our experiments, we also measure the fraction of instances successfully solved without this issue.
Experiment Results and Discussion
Our results on the five different distributions for i.i.d. random disutilities are shown in Fig. 3.
@@
Figure 3: Numerical results on iid randomly generated disutility matrix instances. Each row corresponds to a different disutility distribution (noted on the left). Left: Average number of iterations to solve a given instance size. Middle: Average wall clock running time to solve a given instance size (using Gurobi for both algorithms). Right: Fraction of instances solved for each algorithm. GFW solves every instance for every distribution.
@@
Overall, we conclude that GFW is the first highly practical algorithm: it solves every instance in very little time and in few iterations, even for instances where the size of the disutility matrix is $300\times 300$. In comparison, EPM fails to solve many instances in our datasets. As discussed before, we do caution that the success of EPM is highly dependent on the performance of the Gurobi QP solver, and in particular its accuracy. We found that the QP performance of Gurobi can vary substantially across different machines, and unpredictably so. Due to this so-called “performance variability”, it is difficult to conclusively evaluate the wall-clock performance of the algorithms, as both algorithms rely on the performance of Gurobi on large-size instances. In particular, some preliminary experiments suggests that the wall-clock performance of EPM improves relative to the wall-clock performance of GFW if one adds more RAM. However, we found that even in this case, EPM still has a significant failure rate, since Gurobi still outputs approximate QP solutions, and this still leads to numerical errors that break EPM.
@@
Figure 4: Numerical results on AAMAS PC member bidding data. The top row uses the “original” valuations. The bottom row adds Gaussian noise. Left: Average number of iterations to solve a given instance size. Middle: Average wall clock running time to solve a given instance size. Right: Fraction of instances solved.
6 Conclusion and Discussion
We develop a new convex maximization program for computing CE in Fisher markets with chores, and showed that it satisfies some dual-like properties with the concave minimization program for computing CE. We then derived a new redundant constraint which allows us to avoid directions that lead to infinitely positive objective. This yielded the first mathematical program for finding CE for chores such that it is possible to run iterative optimization methods over a convex feasible region. We then introduced the GFW method for computing CE. From prior results on convex maximization over polyhedra, it follows that this method terminates in a finite number of iterations. We then showed that, in fact, the method finds an $\epsilon$-approximate CE in $O(\frac{1}{\epsilon^{2}})$ iterations, while requiring solving only a simple LP at every iteration. We also gave a more general convergence result that generalizes prior results for strongly convex functions to Bregman-strongly-convex functions. Finally, we showed through extensive numerical experiments that our method is highly practical: it solved every instance that we tried in a minute or less and is both robust and simple to implement. In contrast, we showed that the prior state-of-the-art method, the EPM, often fails to solve instances altogether, due to numerical accuracy problems that break the algorithm.
Challenge of moving from approximate CE to exact CE.
The obvious open problem is to settle the exact complexity of finding a CE in the chores market. A natural attempt would be to find a good approximate CE and then argue how to move to a “nearby” exact CE. Unfortunately, such approaches can face the issue that there are instances where an approximate CE can be very far from an exact CE (See Appendix B for an example).
@@ 6 Conclusion and Discussion
Despite the above hurdle, we believe that finding a $\epsilon$-approximate CE in the chores market with a better time complexity dependence on $\frac{1}{\epsilon}$ is of theoretical interest. An ideal goal would be to have a poly-logarithmic dependence on $\frac{1}{\epsilon}$, but any polynomial improvement would be a stepping stone (currently, all methods have a dependence of $\mathcal{O}(\frac{1}{\epsilon^{2}})$).
Acknowledgement
Christian Kroer was supported by the Office of Naval Research awards N00014-22-1-2530 and N00014-23-1-2374, and the National Science Foundation awards IIS-2147361 and IIS-2238960. Ruta Mehta was supported by the NSF grant CCF-2334461.
@@ Appendix A A Primal-Dual Analysis of the Convergence of the Frank Wolfe Method to an Approximate Equilibrium
Since $f$ does not depend on $p$, we ease notation in this by section by writing $f(\beta)$ for the (Chores Dual) objective function, which was previously defined as depending on both $p$ and $\beta$.
@@
$\min_{\beta}$ $\sum_{i\in[n]}B_{i}\cdot\frac{\beta_{i}}{\beta^{t-1}_{i}}$ (19)
@@
Let $x^{t}_{ij}$ be the optimal dual variable corresponding to the constraint $p_{j}-\beta_{i}d_{ij}\leq 0$, and $\delta^{t}$ be the optimal dual variable corresponding to the constraint $\sum_{j\in[m]}p_{j}=1$ in the $t$-th iteration of the greedy Frank Wolfe algorithm.
Claim 2.
@@
Under stationarity conditions of the optimum solutions, we have for all $i\in[n]$, $B_{i}/\beta^{t-1}_{i}=\sum_{j\in[m]}d_{ij}x^{t}_{ij}$, whenever $\beta^{t}_{i}>0$. Thus, it suffices to show that $\beta^{t}_{i}>0$ for all $i\in[n]$. Assume otherwise, and $\beta^{t}_{\ell}=0$. Since $\beta^{t}_{\ell}=\max_{j\in[m]}p^{t}_{j}/d_{\ell j}$, this implies that $p^{t}_{j}=0$ for all $j\in[m]$, implying $\sum_{j\in[m]}p^{t}_{j}=0$, which is a contradiction. ∎
@@
From 3, it follows that $\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t-1}_{ij})^{B_{i}}=(\prod_{i\in[n]}(1/\beta^{t-1}_{i})^{B_{i}})\cdot(\prod_{i\in[n]}(\sum_{j\in[m]}$ $p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}})$ and 2 it follows that $\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t}_{ij})^{B_{i}}=\prod_{i\in[n]}(B_{i}/\beta^{t-1}_{i})^{B_{i}}$ Therefore, if we comparing the weighted product of disutilities in the current iteration and the previous iteration, we have
$\frac{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t}_{ij})^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}d_{ij}x^{t-1}_{ij})^{B_{i}}}$
$=$ $\frac{\prod_{i\in[n]}(\frac{B_{i}}{\beta^{t-1}_{i}})^{B_{i}}}{(\prod_{i\in[n]}(\frac{1}{\beta^{t-1}_{i}})^{B_{i}})\cdot(\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}})}$
@@
We now make an observation on $\sum_{i\in[n]}x^{t}_{ij}$– interpreting the dual variables $x^{t}_{ij}$ as the fraction of chore $j$ allocated to agent $i$, one can interpret $\sum_{i\in[n]}x^{t}_{ij}$ as the total consumption of chore $j$ at the end of the $t$-th iteration.
Claim 4.
@@
Under stationarity conditions, we have $\sum_{i\in[n]}x^{t}_{ij}=-\delta^{t}$ for all $j$ such that $p^{t}_{j}>0$. Furthermore, from 3, it follows that $\sum_{j\in[m]}p^{t}_{j}x^{t}_{ij}=\beta^{t}_{i}\cdot(\sum_{j\in[m]}d_{ij}x^{t}_{ij})$. Observe that,
@@
$=\sum_{i\in[n]}\beta^{t}_{i}\cdot(\sum_{j\in[m]}d_{ij}x^{t}_{ij})$ (by 2)
$=\sum_{i\in[n]}\sum_{j\in[m]}p^{t}_{j}x^{t}_{ij}$ (by 3)
@@
$=\sum_{j\mid p^{t}_{j}>0}p^{t}_{j}\cdot(-\delta^{t})$
$=(-\delta^{t})\cdot\sum_{j\in[m]}p^{t}_{j}$
@@
We prove this by contradiction. Assume there exists a $j\in[m]$, such that $p^{t}_{j}=0$. Then note that $p^{t}_{j}/d_{ij}<\beta^{t}_{i}$ for all $i\in[n]$ (follows from the fact that every $\beta^{t}_{i}$ is strictly positive). Note that there exists a $\gamma\ll\min_{\ell\in[n]}d_{\ell j}\beta^{t}_{\ell}$, such that if we increase $p^{t}_{j}$ to $\gamma$, and decrease all other prices by a multiplicative factor of $(B-\gamma)$, we still have a feasible set of prices [^9] All the prices are still non-negative and they still sum up to $B$. and $\max_{j\in[m]}p^{t}_{j}/d_{ij}$ strictly decreases for all $i\in[n]$. Let $\beta^{\prime}_{i}=\max_{j\in[m]}p^{t}_{j}/d_{ij}$ (post the multiplicative scaling). Note that $\beta^{\prime}_{i}<\beta^{t}_{i}$ for all $i\in[n]$ and thus $\sum_{i\in[n]}B_{i}\cdot\beta^{\prime}_{i}/\beta^{t-1}_{i}<\sum_{i\in[n]}B_{i}\cdot\beta^{t}_{i}/\beta^{t-1}_{i}$ which contradicts the optimality of $\beta^{t}$. ∎
Claim 5.
If $(p^{t-1},x^{t-1})$ is not an $\epsilon$-approximate equilibrium, then $\log\Bigg{(}\frac{\prod_{i\in[n]}B_{i}^{B_{i}}}{\prod_{i\in[n]}(\sum_{j\in[m]}p^{t-1}_{j}x^{t-1}_{ij})^{B_{i}}}\Bigg{)}\geq\min_{i}B_{i}\cdot\epsilon^{2}/2.3$.